# {textblock}

### v1

@@ Abstract
We present AIOS 1.0, a novel platform designed to advance computer-use agent (CUA) capabilities through environmental contextualization. While existing approaches primarily focus on building more powerful agent frameworks or enhancing agent models, we identify a fundamental limitation: the semantic disconnect between how language models understand the world and how computer interfaces are structured. AIOS 1.0 addresses this challenge by transforming computers into contextual environments that language models can natively comprehend, implementing a Model Context Protocol (MCP) server architecture to abstract computer states and actions. This approach effectively decouples interface complexity from decision complexity, enabling agents to reason more effectively about computing environments. To demonstrate our platform’s effectiveness, we introduce LiteCUA, a lightweight computer-use agent built on AIOS 1.0 that achieves a 14.66% success rate on the OSWorld benchmark, outperforming several specialized agent frameworks despite its simple architecture. Our results suggest that contextualizing computer environments for language models represents a promising direction for developing more capable computer-use agents and advancing toward AGI that can interact with digital systems. The source code of LiteCUA is available at GitHub1 and it is also integrated into the AIOS main branch as part of AIOS2.
[Uncaptioned image] 1LiteCUA Code Repository: https://github.com/agiresearch/LiteCUA
[Uncaptioned image] 2AIOS Code Repository: https://github.com/agiresearch/AIOS
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023, Team et al., 2023, Liu et al., 2024, Dubey et al., 2024, Yang et al., 2024a) have achieved impressive leaps in understanding knowledge and reasoning (Wei et al., 2022, Yao et al., 2023a, Besta et al., 2023, Jin et al., 2024b, 2025, Chen et al., 2022, Jin et al., 2024a), fuelling a new generation of LLM-powered agents that plan, observe, and act to interact with digital environments (Shen et al., 2023, Wu et al., 2023, Ge et al., 2023, Yao et al., 2023b). Research demonstrates that, when coupled with suitable prompting or fine-tuning, LLM-based agents can effectively plan to fulfill complex tool calling (Patil et al., 2024, Qin et al., 2024, Xu et al., 2023) and programming (Wang et al., 2024, Yang et al., , Zhang et al., 2024) tasks.
Moving beyond calling external tools or executing generated code, researchers have begun exploring whether LLM agents can perform more complex computer-use tasks, leveraging their emerging planning and action capabilities. This exploration has catalyzed the development of benchmarks assessing LLM-based agents in computer-use scenarios across operating systems (Xie et al., 2024, Bonatti et al., 2024), mobile devices (Deng et al., 2024, Rawles et al., 2024), and browsers (Yao et al., 2022, Zhou et al., 2023, Deng et al., 2023, Koh et al., 2024, Pan et al., 2024). Parallel efforts have focused on leveraging LLMs and vision-language models (VLMs) to perceive computer/mobile environments and execute actions such as clicking, typing, and dragging interactive elements to complete tasks. Current approaches to computer-use agents (CUAs) primarily follow two trajectories: (1) developing agent frameworks with specialized modules for critical subtasks like grounding (Agashe et al., 2025, Gou et al., 2025), search (Yu et al., 2024), and reflection (Zhang et al., 2025b, Wang and Liu, 2024); and (2) embedding agentic capabilities directly into LLMs/VLMs to create native agent models (Qin et al., 2025, Guo et al., 2025). Despite these advances, CUAs continue to exhibit significant limitations in their ability to navigate complex computing environments effectively.
@@
Inspired by this and based on the foundation laid out by the AIOS architecture (Mei et al., 2024, Rama et al., 2025), we introduce AIOS 1.0, an augmented platform developed to move towards contextualization of computer environments for LLM-based agents. Specifically, AIOS 1.0 aims to model the full computer as MCP [^1] https://modelcontextprotocol.io/introduction server to provide LLM-friendly and well-structured context where interactive elements permissible actions are explicitly encoded in JSON schemas. This contextualization process transforms the computer from an opaque operational environment into a semantic landscape that aligns with the LLM’s understanding of the environment. This design equips agents with an informative yet compact view of the state space, enabling finer-grained reasoning over available operations and facilitating longer-horizon planning. By decoupling interface complexity from decision complexity, AIOS 1.0 addresses a fundamental bottleneck in current CUA capabilities. The contextualized environment acts as a semantic bridge between the agent’s language-based reasoning and the computer’s operational requirements.
Upon AIOS 1.0, we introduce LiteCUA, a computer-use agent that is developed by us with simple architectural design. LiteCUA achieves competitive results with several agent frameworks, which scores 14.66% in the OSWorld benchmark (Xie et al., 2024). The success of LiteCUA emphasizes that AIOS 1.0’s approach of creating agent-specific environments can effectively reduce the implementation complexity required to build computer-use agents, allowing researchers to focus on advancing agent capabilities rather than wrestling with the intricacies of computer interface interpretation. By reframing the computer-use agent challenge as an environmental contextualization problem rather than solely a model capability issue, AIOS 1.0 aims to catalyze the next leap in computer-use agent capabilities. This approach represents a crucial step toward developing truly general computer-use agents, systems that can navigate the digital world with the flexibility and comprehension previously exclusive to human users.
@@ 2 Architecture and System Design
Following the architecture of AIOS (Mei et al., 2024, Rama et al., 2025), which abstracts a specific AIOS kernel for isolating agent logic (Rama et al., 2025) and managing agent-related resources and services, we extend this architecture to create AIOS 1.0, a platform optimized for developing and running computer-use agents.
Figure 2: The serving architecture of AIOS 1.0 for computer-use agent, which extends on the basis of AIOS 0.x versions (Mei et al., 2024).
@@ 2.1 Architecture
Application Layer: In the Application Layer, agents are provided with SDK APIs to interact with some crucial components like Terminal, Code editor, Browser, and Document. By abstracting the essential interaction patterns common across all interfaces, AIOS 1.0 creates a consistent interaction surface that allows agents to navigate diverse computing environments. This approach significantly reduces the complexity agents face when operating in heterogeneous computer components. The agents can leverage the system-level resources, such as LLM and vector database by revoking the APIs within AIOS SDK (Rama et al., 2025). While maintaining the original design from AIOS SDK regarding LLM, Memory and Storage Query APIs, AIOS 1.0 designs the integration of MCP for tool communications. MCP Client and Http Client are established as communication channels to map the agent’s semantic understanding into computer manipulation commands.
@@ 3.1 Orchestrator-Worker Architecture
As illustrated in Figure 4, LiteCUA adopts an orchestrator-worker architecture that distributes cognitive responsibilities across specialized modules while maintaining cohesive task execution, which is inspired by some popular agent frameworks (Wu et al., 2023, Agashe et al., 2025, Li et al., 2023).
@@ 4.1 Setup
We evaluate the LiteCUA computer-use agent built atop of AIOS 1.0 on OSWorld (Xie et al., 2024), a benchmark which provides a diverse online environment for evaluating multi-modal agents on computer-using tasks. It consists of 369 tasks involving real-world web and desktop applications.
@@
For baselines, we choose three LLM baselines, i.e., GPT (4o-mini, GPT-4o) (Achiam et al., 2023), and Gemini-1.5-flash (Team et al., 2023), and three agent baselines which adopt GPT-4o as their backbones, i.e., Friday (Wu et al., 2024), Open-Interpreter (Lucas, 2024), and AgentStore (Jia et al., 2024). For a fair comparison, we evaluate all the methods under the "Screenshot + A11y tree" setting and only compare agents which do not introduce extra models for grounding. The metric we use is success rate provided within OSWorld.
@@ 4.2 Comparative Performance
GPT-4o-mini (Achiam et al., 2023) 6.21
GPT-4o (Achiam et al., 2023) 11.21
Gemini-1.5-pro (Team et al., 2023) 5.1
@@
Friday (Wu et al., 2024) 11.11
Open-Interpreter (Lucas, 2024) 8.94
AgentStore (Jia et al., 2024) 13.55
@@ 5.1 Computer-Use Agent
The automation of computer-based tasks through agents has significantly advanced with (M)LLMs, enabling them to operate across desktop, mobile, and web platforms. A key technical improvement involves moving beyond reliance on potentially noisy textual representations like HTML or accessibility trees towards more robust, vision-centric perception. For instance, systems like UGround (Gou et al., 2025) and Aria-UI (Yang et al., 2024b) champion agents that perceive GUIs visually, akin to humans, and perform pixel-level operations. UGround leverages large-scale synthetic web data and adapts the LLaVA architecture for universal visual grounding , while Aria-UI employs a pure-vision approach with a scalable data pipeline to synthesize diverse instruction samples for grounding and incorporates action histories for context-aware reasoning. On mobile platforms, AutoDroid demonstrates a technical advancement by combining LLM commonsense with app-specific knowledge, acquired through automated dynamic analysis and exploration-based memory injection, to automate tasks without prior manual effort (Wen et al., 2024). The evaluation of such computer-use agents is also evolving, with benchmarks like AgentBench (Liu et al., 2023), TheAgentCompany (Xu et al., 2024), and AgentGym (Xi et al., 2023) providing more complex and realistic scenarios for assessing their operational capabilities. Furthermore, the convergence of API-based and GUI-based interaction, creating hybrid agents, is a promising direction for versatile software automation (Qiao et al., 2025).
@@ 5.2 AI Agent Operating System
The increasing complexity of LLM-based agents calls for dedicated infrastructure, leading to the development of AI Agent Operating Systems designed to provide foundational services. AIOS (Mei et al., 2024), as the first LLM agent operating system, introduces an LLM-specific kernel to manage resources and LLM-specific services, offering agent scheduling, context management, memory and storage management, tool management, and access control, along with the agent SDK (Rama et al., 2025) for agent development. Such systems address critical challenges like efficient resource allocation, a problem also explored by research into self-resource allocation in multi-agent LLM systems where LLMs themselves can act as planners for task distribution (Amayuelas et al., 2025). UFO2 (Zhang et al., 2025a) designs and implements the desktop agent OS within the Windows environment. These developments show that agent-specific environments are crucial for creating scalable, reliable, and secure ecosystems for advanced agent applications.

### v2

# LiteCUA: Computer as MCP Server for Computer-Use Agent on AIOS

@@ Abstract
We present AIOS 1.0, a novel platform designed to advance computer-use agent (CUA) capabilities through environmental contextualization. While existing approaches primarily focus on building more powerful agent frameworks or enhancing agent models, we identify a fundamental limitation: the semantic disconnect between how language models understand the world and how computer interfaces are structured. AIOS 1.0 addresses this challenge by transforming computers into contextual environments that language models can natively comprehend, implementing a Model Context Protocol (MCP) server architecture to abstract computer states and actions. This approach effectively decouples interface complexity from decision complexity, enabling agents to reason more effectively about computing environments. To demonstrate our platform’s effectiveness, we introduce LiteCUA, a lightweight computer-use agent built on AIOS 1.0 that achieves a 14.66% success rate on the OSWorld benchmark, outperforming several specialized agent frameworks despite its simple architecture. Our results suggest that contextualizing computer environments for language models represents a promising direction for developing more capable computer-use agents and advancing toward AGI that can interact with digital systems.
@@ 1 Introduction
Large language models (LLMs) (OpenAIGPT4, ; team2023gemini, ; liu2024deepseek, ; dubey2024llama, ; yang2024qwen2, ) have achieved impressive leaps in understanding knowledge and reasoning (wei2023chainofthought, ; yao2023tree, ; besta2023graph, ; jin2024impact, ; jin2025two, ; chen2022program, ; jin2024learning, ), fuelling a new generation of LLM-powered agents that plan, observe, and act to interact with digital environments (shen2023hugginggpt, ; wu2023autogen, ; ge2023openagi, ; yao2023react, ). Research demonstrates that, when coupled with suitable prompting or fine-tuning, LLM-based agents can effectively plan to fulfill complex tool calling (patil2024gorilla, ; qin2023toolllm, ; xu2023tool, ) and programming (wang2024executable, ; yangswe, ; zhang2024codeagent, ) tasks.
Moving beyond calling external tools or executing generated code, researchers have begun exploring whether LLM agents can perform more complex computer-use tasks, leveraging their emerging planning and action capabilities. This exploration has catalyzed the development of benchmarks assessing LLM-based agents in computer-use scenarios across operating systems (xie2024osworld, ; bonatti2024windows, ), mobile devices (deng2024mobile, ; rawles2024androidworld, ), and browsers (yao2022webshop, ; zhou2023webarena, ; deng2023mind2web, ; koh2024visualwebarena, ; pan2024webcanvas, ). Parallel efforts have focused on leveraging LLMs and vision-language models (VLMs) to perceive computer/mobile environments and execute actions such as clicking, typing, and dragging interactive elements to complete tasks. Current approaches to computer-use agents (CUAs) primarily follow two trajectories: (1) developing agent frameworks with specialized modules for critical subtasks like grounding (agashe2025agent, ; gou2025navigating, ), search (yu2024exact, ), and reflection (zhang2025enhancing, ; wang2024oscar, ); and (2) embedding agentic capabilities directly into LLMs/VLMs to create native agent models (qin2025ui, ; guo2025seed1, ). Despite these advances, CUAs continue to exhibit significant limitations in their ability to navigate complex computing environments effectively.
@@
Inspired by this and based on the foundation laid out by the AIOS architecture (mei2024aios, ; rama2025cerebrum, ), we introduce AIOS 1.0, an augmented platform developed to move towards contextualization of computer environments for LLM-based agents. Specifically, AIOS 1.0 aims to model the full computer as MCP [^1] https://modelcontextprotocol.io/introduction server to provide LLM-friendly and well-structured context where interactive elements permissible actions are explicitly encoded in JSON schemas. This contextualization process transforms the computer from an opaque operational environment into a semantic landscape that aligns with the LLM’s understanding of the environment. This design equips agents with an informative yet compact view of the state space, enabling finer-grained reasoning over available operations and facilitating longer-horizon planning. By decoupling interface complexity from decision complexity, AIOS 1.0 addresses a fundamental bottleneck in current CUA capabilities. The contextualized environment acts as a semantic bridge between the agent’s language-based reasoning and the computer’s operational requirements.
Upon AIOS 1.0, we introduce LiteCUA, a computer-use agent that is developed by us with simple architectural design. LiteCUA achieves competitive results with several agent frameworks, which scores 14.66% in the OSWorld benchmark (xie2024osworld, ). The success of LiteCUA emphasizes that AIOS 1.0’s approach of creating agent-specific environments can effectively reduce the implementation complexity required to build computer-use agents, allowing researchers to focus on advancing agent capabilities rather than wrestling with the intricacies of computer interface interpretation. By reframing the computer-use agent challenge as an environmental contextualization problem rather than solely a model capability issue, AIOS 1.0 aims to catalyze the next leap in computer-use agent capabilities. This approach represents a crucial step toward developing truly general computer-use agents, systems that can navigate the digital world with the flexibility and comprehension previously exclusive to human users.
@@ 2 Architecture and System Design
Following the architecture of AIOS (mei2024aios, ; rama2025cerebrum, ), which abstracts a specific AIOS kernel for isolating agent logic (rama2025cerebrum, ) and managing agent-related resources and services, we extend this architecture to create AIOS 1.0, a platform optimized for developing and running computer-use agents.
Figure 2: The serving architecture of AIOS 1.0 for computer-use agent, which extends on the basis of AIOS 0.x versions (mei2024aios, ).
@@ 2.1 Architecture
Application Layer: In the Application Layer, agents are provided with SDK APIs to interact with some crucial components like Terminal, Code editor, Browser, and Document. By abstracting the essential interaction patterns common across all interfaces, AIOS 1.0 creates a consistent interaction surface that allows agents to navigate diverse computing environments. This approach significantly reduces the complexity agents face when operating in heterogeneous computer components. The agents can leverage the system-level resources, such as LLM and vector database by revoking the APIs within AIOS SDK (rama2025cerebrum, ). While maintaining the original design from AIOS SDK regarding LLM, Memory and Storage Query APIs, AIOS 1.0 designs the integration of MCP for tool communications. MCP Client and Http Client are established as communication channels to map the agent’s semantic understanding into computer manipulation commands.
@@ 3.1 Orchestrator-Worker Architecture
As illustrated in Figure 4, LiteCUA adopts an orchestrator-worker architecture that distributes cognitive responsibilities across specialized modules while maintaining cohesive task execution, which is inspired by some popular agent frameworks (wu2023autogen, ; agashe2025agent, ; li2023camel, ).
@@ 4.1 Setup
We evaluate the LiteCUA computer-use agent built atop of AIOS 1.0 on OSWorld (xie2024osworld, ), a benchmark which provides a diverse online environment for evaluating multi-modal agents on computer-using tasks. It consists of 369 tasks involving real-world web and desktop applications.
@@
For baselines, we choose three LLM baselines, i.e., GPT (4o-mini, GPT-4o) (OpenAIGPT4, ), and Gemini-1.5-flash (team2023gemini, ), and three agent baselines which adopt GPT-4o as their backbones, i.e., Friday (wu2024copilot, ), Open-Interpreter (interpreter2024, ), and AgentStore (jia2024agentstore, ). For a fair comparison, we evaluate all the methods under the "Screenshot + A11y tree" setting and only compare agents which do not introduce extra models for grounding. The metric we use is success rate provided within OSWorld.
@@ 4.2 Comparative Performance
GPT-4o-mini (OpenAIGPT4, ) 6.21
GPT-4o (OpenAIGPT4, ) 11.21
Gemini-1.5-pro (team2023gemini, ) 5.1
@@
Friday (wu2024copilot, ) 11.11
Open-Interpreter (interpreter2024, ) 8.94
AgentStore (jia2024agentstore, ) 13.55
@@ 5.1 Computer-Use Agent
The automation of computer-based tasks through agents has significantly advanced with (M)LLMs, enabling them to operate across desktop, mobile, and web platforms. A key technical improvement involves moving beyond reliance on potentially noisy textual representations like HTML or accessibility trees towards more robust, vision-centric perception. For instance, systems like UGround (gou2025navigating, ) and Aria-UI (yang2024ariaui, ) champion agents that perceive GUIs visually, akin to humans, and perform pixel-level operations. UGround leverages large-scale synthetic web data and adapts the LLaVA architecture for universal visual grounding , while Aria-UI employs a pure-vision approach with a scalable data pipeline to synthesize diverse instruction samples for grounding and incorporates action histories for context-aware reasoning. On mobile platforms, AutoDroid demonstrates a technical advancement by combining LLM commonsense with app-specific knowledge, acquired through automated dynamic analysis and exploration-based memory injection, to automate tasks without prior manual effort (wen2023autodroid, ). The evaluation of such computer-use agents is also evolving, with benchmarks like AgentBench (liu2023agentbench, ), TheAgentCompany (xu2024agentcompany, ), and AgentGym (xi2023agentgym, ) providing more complex and realistic scenarios for assessing their operational capabilities. Furthermore, the convergence of API-based and GUI-based interaction, creating hybrid agents, is a promising direction for versatile software automation (qiao2025apivsgui, ).
@@ 5.2 AI Agent Operating System
The increasing complexity of LLM-based agents calls for dedicated infrastructure, leading to the development of AI Agent Operating Systems designed to provide foundational services. AIOS (mei2024aios, ), as the first LLM agent operating system, introduces an LLM-specific kernel to manage resources and LLM-specific services, offering agent scheduling, context management, memory and storage management, tool management, and access control, along with the agent SDK (rama2025cerebrum, ) for agent development. Such systems address critical challenges like efficient resource allocation, a problem also explored by research into self-resource allocation in multi-agent LLM systems where LLMs themselves can act as planners for task distribution (amayuelas2025selfresource, ). UFO2 (zhang2025ufo2, ) designs and implements the desktop agent OS within the Windows environment. These developments show that agent-specific environments are crucial for creating scalable, reliable, and secure ecosystems for advanced agent applications.